OpenStudy (haleyelizabeth2017):

Find all the zeros of the equation. I know 1 is one of the zeros, but I can't remember how to find the rest :(

OpenStudy (haleyelizabeth2017):

\[3x^2-4=-x^4\]

OpenStudy (anonymous):

\[\begin{align*}3x^2-4&=-x^4\\ x^4+3x^2-4&=0\\ z^2+3z-4&=0&\text{where }z=x^2 \end{align*}\] Factor from here, or use the quadratic formula.

OpenStudy (haleyelizabeth2017):

I'm confused on how you got z....

OpenStudy (anonymous):

I implemented a substitution. It's not essential to solving the equation, but it helps you to approach the problem as something much simpler. Observe that \[\underbrace{x^4+3x^2-4}_{(1)}=(x^2)^2+3x^2-4\] In other words, the expression is quadratic in \(x^2\), so if I let \(z=x^2\), then naturally the above is the same as \[\underbrace{z^2+3z-4}_{(2)}\] You can factor this to get \[\underbrace{(z+4)(z-1)}_{(3)}\] which is equivalent to \[\underbrace{(x^2-4)(x^2-1)}_{(4)}\] There's nothing stopping you from jumping from expression (1) to (4), but the intermediate steps (2) and (3) sort of help to demonstrate that this problem can be solved as a quadratic equation.

OpenStudy (haleyelizabeth2017):

Okay.

OpenStudy (haleyelizabeth2017):

So \(27x^2-324=-x^4\) would pretty much be solved the same way?